# Kekulé order from diffuse nesting near higher-order Van Hove points

### v1

@@ A Breathing kagome model
Figure SM 1: Kinetic model for the breathing kagome lattice. The hopping processes $t_{1,2,3,4}$ from sublattice $A$ to the neighbouring unit cells are shown with circles illustrating the hopping distance. The hopping $t_{4}$ vanishes along a line perpendicular to the residual mirror symmetry at each sublattice (purple). In Co3Sn2S2, the blue sites correspond to Co atoms and the yellow sites to S atoms which hybridise with the kagome lattice of Co, driving the breathing distortion.
The kinetic model reduces to the ordinary nearest-neighbor kagome model in the case of $t_{1}=t_{2}$, $t_{3}=t_{4}=0$, which exhibits an ordinary VHS at the $M$-point. The HOVHS seen in DFT is captured by the choice $t_{1\rightarrow 4}\approx(-1.383,-0.617,0.352,-0.176)\,t$ with the the overall scale $t\approx 0.09$ eV [71]. This leads to a vanishing second derivative of the energy dispersion along the $y$-direction at $M_{2}=\tfrac{2\pi}{a\sqrt{3}}(0,1)$. In particular, it is possible to show that the second derivative vanishes along a continuous line through $t_{3}=-2t_{4}$ in the $t_{3}-t_{4}$ parameter space when keeping $(t_{1},t_{2})=\frac{2}{\frac{161}{41}+\frac{7}{4}}(-\tfrac{161}{41},-\tfrac{7}{4})t$ fixed [71]. The energy dispersion in the lowest band is described by $\varepsilon(M_{2}+\bm{k})=\alpha k_{x}^{2}-\beta k_{y}^{4}$ with positive constants $\alpha,\beta$. The resulting HOVHS can be classified as belonging to the $A_{3}$ class [127, 79].
@@ B Non-breathing kagome model
It is possible to achieve the same type of HOVHS even without the breathing distortion in our model, i.e., with $t_{1}=t_{2}$. One possible choice is $(t_{1},t_{2},t_{3},t_{4})=(-1,-1,0.370845,-0.127329)\,t$. This restores the full $C_{6v}$ symmetry of the original kagome lattice. In analogy to the breathing scenario, a continuous line is present in the $t_{3}-t_{4}$ parameter space preserving the vanishing of the quadratic part.
@@
C Lattice regularized low energy effective model
To track the origin of diffuse nesting to the Van Hove singularities, we employ a Taylor approximated version of the above model. To this end, we first project the model to its lowest band that contains the HOVHS. This single-band energy dispersion is expanded around the M points up to 4th order, which captures all leading contributions. To find the canonical dispersion [127], i.e. the leading scale-invariant part of the dispersion defined by the quasi-homogeneity condition $E(\lambda^{a}k_{x},\lambda^{b}k_{y})=\lambda E(k_{x},k_{y})$, we identify the scaling exponents of the HOVHS as $a=\frac{1}{2},b=\frac{1}{4}$. This leads to a canonical dispersion of the form $E(k_{x},k_{y})=Ak_{x}^{2}+Bk_{x}k_{y}^{2}+Ck_{y}^{4}$. The prefactors of the HOVHS in the breathing Kagome model are $A=0.258714795,B=0,C=-8.63917358\cdot 10^{-3}$ when the axes are chosen such that the $k_{x}$ direction is on the $MK$ line and the $k_{y}$ direction on the $M\Gamma$ line. To define the energy dispersion of the Taylor model on the complete hexagonal Brillouin zone, we divide the latter into six equilateral triangles. Each of them has a single M point on its boundary. We define the energy dispersion as a piecewise function on these triangles, where in each triangle the canonical dispersion of the corresponding M point is used. This approach conserves the lattice symmetry and ensures that the dispersion is continuous. It is, however, not differentiable on the $\Gamma K$ lines.
@@ C Lattice regularized low energy effective model
We employ the functional renormalization group in its standard static four-point approximation, neglecting self-energy corrections, frequency dependencies, and all interaction vertices with more than 4 fermionic fields, i.e. two interacting particles [103, 104, 105, 106, 107]. The main methodological improvement compared to previous studies of HOVHS [86, 87, 111, 112, 113] is that within the aforementioned approximations our FRG analysis systematically accounts for all possible ordering tendencies and fluctuations representable as fermionic bilinears with arbitrary momentum transfers $\boldsymbol{q}$ in an unbiased fashion. While secondary momenta $\boldsymbol{k}$ and $\boldsymbol{k}^{\prime}$ are expanded in real-space form factors that are truncated at some distance $\alpha$ (truncated unity FRG formulation), transfer momenta are sampled on equidistant momentum points within the Brillouin zone, where the modern implementation used in the present work [106, 107, 108, 109] allows for high resolutions up to $\approx 25000$ momentum mesh points within the BZ.
@@
Since the FRG is only valid directly at the phase transition—where the divergence is encountered—the mean-field decomposition is performed at the phase boundary where $\Delta\rightarrow 0$. Consequently, we can linearize the self-consistent gap equation, recasting it as an eigenvalue problem for each $\boldsymbol{q}$ directly at the phase transition. The maximum of eigenvalues at each transfer momentum in the particle-hole channel of our model, indicating the leading fluctuation, is displayed on an exemplary momentum mesh in Fig. 2a. The associated eigenstate of a divergent eigenvalue defines the structure of the leading order parameter.
@@ A Interaction parameter phase diagrams
To investigate the ordering tendencies for the models introduced in Section SM 1, we add density-density interactions to the tight-binding Hamiltonian up to fourth nearest neighbor distance,
@@
The predominant part of parameter space of the breathing model consists of the unprecedented $\sqrt{3}\times\sqrt{3}$ density wave order originating from the diffuse nesting mechanism introduced by the present work. In the limit of weak initial interactions, superconductivity and nematic orders emerge as the dominant tendencies. The latter is in agreement with previous theoretical and experimental studies of Co3Sn2S2 [71] and theoretical studies of HOVHS scenarios in hexagonal lattice systems [86, 87]. Upon introducing a spin degree of freedom, the leading instabilities reveals a magnetic ordering tendency rather than charge order; the alternating bond phase then corresponds to the spins rather than charge density. While we comprehensively analyzed the mean field equations and Ginzburg-Landau free energy for the scenario of spinless fermions, we leave a comparably thorough study of magnetic Kekulé order to future works.
@@ B $K$-order stability analysis
To extend the stability analysis of the $K$-ordered phase beyond the interaction parameter regimes, we vary multiple parameters of the kinetic model given in Eq. 1 and repeat the FRG analysis.
Deviations from van Hove filling
We observe a suppression of the $K$-ordering tendency upon small doping. While the order persists within the regime of intermediate initial interaction strengths for dopings $\mu\leq 0.05$, it is absent within the interaction parameter range accessible to the FRG calculation for $\mu\geq 0.1$.
Deviations from the higher-order van Hove scenario
As outlined earlier, a van Hove singularity is classified as higher-order if the quartic terms of the dispersion are finite while the quadratic terms vanish. We analyze the effect of small deviations from this scenario by varying the hybridization parameters of the model presented in Eq. 1, thereby reintroducing a finite quadratic term along the $\Gamma M$ line. At sufficiently flat quadratic dispersion, the $K$ order persists, demonstrating that the flatness-induced high DOS of the VHS close to its higher-order configuration drives the mechanism while an exact vanishing of the Hessian determinant is not required. As the deviation is increased, we observe that the transfer momentum associated with the ordering tendency shifts toward an incommensurate momentum point located along the $KM$ line. We find that this is consistent with a shift in the momentum of the particle-hole susceptibility peak, further substantiating the claim that our effect is purely of kinetic origin. However, if longer-range interactions that favor the charge imbalance between neighboring up-pointing triangles are strong, the $K$-order can be stabilized even though the particle-hole susceptibility peak shifts to incommensurate momenta.
@@
The particle-hole susceptibility peak at transfer-momentum $\boldsymbol{q}=K$ in Fig. 3 (main text) is traced back to the diffuse Fermi surface (FS) nesting inherent to the warped FS emerging in the precence of HOVHSs, where the quartic term points along the $\Gamma M$ direction. The Fermi surface overlaps at $\boldsymbol{q}=K$ and $\boldsymbol{q}=M$, with broadening due to temperature $T=0.01t$, are shown as heatmaps in Fig. SM 4. The gray regions indicate the two shifted FS’s. In the $K$-shifted scenario extended circular nesting regions appear, while the nesting at $M$-shifting remains poor. This is in contrast to the perfect VHS nesting emerging for ordinary VHS at transfer momenta $M$ in hexagonal lattice systems with only (real) nearest neighbor hybridization.
@@
Figure SM 4: Fermi surface overlaps of breathing and non-breathing kinetic models with nesting-vectors $\boldsymbol{q}=M$ and $\boldsymbol{q}=K$ at HOVHS filling. The broadening factor is $T=0.01\,t$. Upon $K$-shifting circular regions appear, indicating strong nesting. When shifted by $M$, the nesting is poor.
The inherent lattice geometry of the kagome leads to a complete (partial) sublattice polarization at the pure $p$- (mixed $m$-) type VHS at the $M$ points, protected by $\sigma_{v}$ mirror symmetry. In the breathing configuration, this symmetry is explicitly broken on the single-particle level and the FS states feature support on all three sublattices throughout the Fermi surface. Hence, the nesting analysis in Fig. SM 4 does not require a detailed examination of the electronic eigenstates on the FS. This is further corroborated by the breathing configuration inducing a trivial gap in the spectrum, which permits a single-orbital description of the lowest band hosting the HOVHS. This naturally excludes any sublattice polarization effects from emerging in the low-energy description of the system.
Figure SM 5: Fermi surface nesting in the non-breathing scenario. (a) C sublattice weight along the Fermi surface, indicated by color. The sublattice polarization varies smoothly along the FS and is hence present over an extended region around the $M$-point, where a vanishing C weight is enforced by $C_{2}$ symmetry. (b) Original (red) and $K$-shifted (blue) Fermi surface. The opacity indicates again contribution of sublattice $C$ to the FS states. The FS regions with most weight on C only nests with FS segments with little C contribution (indicated by the black rectangle) reproducing the well-known sublattice interference effect also for $K$-nesting. (c) Same as (b) but now with different sublattice contribution on the shifted Fermi surface, indicating the available scattering processes for nearest-neighbor interactions. Now FS states with high occupation on both the original and shifted FS nest to drive a $K$-ordering.
In the fully $C_{6v}$ symmetric non-breathing scenario, the HOVHS is of mixed sublattice type and the FS around the three inequivalent $M$-points becomes depopulated of weight from a specific sublattice (cf. Fig. SM 5(a)). It is worth noting that due to the continuous evolution of the FS states especially in the FS segments with large DOS close to the HOVHS points along the $\sigma_{v}$ mirror plane, sublattice polarization is still strong. This necessitates a sublattice-resolved analysis of the $\boldsymbol{q}=K$ nesting that differs substantially between the identical FS shapes in Fig. SM 4: The overlap region highlighted in Fig. SM 5(b) indicates a dominant FS segment of low sublattice weight, which is nested with regions of high weight of the same sublattice. These states are therefore inaccessible to local scattering events and require a mediation by a NN repulsion that mixes sublattices (compare Fig. SM 5(b) and (c)). This scenario for the $K$-nesting is exactly analogous to the well-known sublattice interference effect for the $M$ point VHS nesting in the kagome Hubbard model [118]. Consequently, we inspect a suppression of critical scales and a disappearing $K$-phase in the small $V_{12}$ regime in Fig. SM 3(c). For large NN repulsion, we recover the usual SC order known from FRG studies on the conventional kagome Hubbard model [115, 6, 114]. While sublattice interference, as originally proposed, favors bond order formation by enhancing longer-range correlations, the simultaneous suppression of the critical scale disfavors the Kekulé bond order, which we have shown to arise from diffuse nesting.

### v2

@@ A Breathing kagome model
Figure SM 1: Kinetic model for the breathing kagome lattice. The hopping processes $t_{1,2,3,4}$ from sublattice AA to the neighbouring unit cells are shown with circles illustrating the hopping distance. The hopping $t_{4}$ vanishes along a line perpendicular to the residual mirror symmetry at each sublattice (purple). In Co3Sn2S2, the blue sites correspond to Co atoms and the yellow sites to S atoms which hybridise with the kagome lattice of Co, driving the breathing distortion.
The kinetic model reduces to the ordinary nearest-neighbor kagome model in the case of $t_{1}=t_{2}$, $t_{3}=t_{4}=0$, which exhibits an ordinary VHS at the $M$-point. In comparing with DFT, a fitting choice reported in Ref. [74] produces a set of hopping parameters $t^{Fit}_{1\rightarrow 4}\approx(-1.5,-0.5,0.312,-0.375)\,t$. We normalize all kinetic parameters such that $t_{1}+t_{2}=2t$. The average of $t_{1}$ and $t_{2}$ is used as energy scale, i.e., $t\approx 0.08\,\mathrm{eV}$. This captures the overall bandstructure and results in a HOVHS at the $M$ points [74]. The bandstructure seen in DFT is in fact closer to the HOVHS scenario than the fitting parameter model; therefore, following Nag et al. [74], we use the mathematically exact HOVHS parameters (with the chemical potential at Van Hove filling) for our theoretical model and reserve the DFT hopping parameters for a stability analysis (cf. Section B) of the diffuse nesting mechanism. An exact HOVHS is captured by the choice $t_{1\rightarrow 4}\approx(-1.383,-0.617,0.352,-0.176)\,t$ with the overall scale $t\approx 0.08$ eV [74]. These parameters lead to a vanishing second derivative of the energy dispersion along the $y$-direction at $M_{2}=\tfrac{2\pi}{a\sqrt{3}}(0,1)$. In particular, it is possible to show that the second derivative vanishes along a continuous line through $t_{3}=-2t_{4}$ in the $t_{3}-t_{4}$ parameter space when keeping $(t_{1},t_{2})=\frac{2}{\frac{161}{41}+\frac{7}{4}}(-\tfrac{161}{41},-\tfrac{7}{4})t$ fixed [74]. The energy dispersion at the Van Hove points is described by $\varepsilon(M_{2}+\bm{k})=\alpha k_{x}^{2}-\beta k_{y}^{4}$ with positive constants $\alpha,\beta$. The resulting HOVHS is classified in Section C as belonging to the $A_{3}$ class [135, 15].
@@ B Non-breathing kagome model
It is possible to achieve the same type of HOVHS even without the breathing distortion in our model, i.e. with $t_{1}=t_{2}$. One possible choice is $(t_{1},t_{2},t_{3},t_{4})=(-1,-1,0.370845,-0.127329)\,t$. This restores the full $C_{6v}$ symmetry of the original kagome lattice. In analogy to the breathing scenario, a continuous line is present in the $t_{3}-t_{4}$ parameter space preserving the vanishing of the quadratic part.
@@
C Low energy effective model
To track the origin of diffuse nesting to the Van Hove singularities, we employ a Taylor approximated version of the above model. To this end, we project the model to its lowest band that contains the HOVHS. We check that this is allowed by investigating the quantum metric of the lowest band, see Section SM 8 for details. The resulting single-band energy dispersion is expanded around the M points up to 4th order, which captures all leading contributions. To find the canonical dispersion [135], i.e., the leading scale-invariant part of the dispersion defined by the quasi-homogeneity condition $E(\lambda^{a}k_{x},\lambda^{b}k_{y})=\lambda E(k_{x},k_{y})$, we identify the scaling exponents of the HOVHS as $a=\frac{1}{2},b=\frac{1}{4}$. This leads to a canonical dispersion of the form $E(k_{x},k_{y})=Ak_{x}^{2}+Bk_{x}k_{y}^{2}+Ck_{y}^{4}$, which corresponds to the $A_{3}$ class [135, 15]. The prefactors of the HOVHS in the breathing Kagome model are (A,B,C)=(0.25871,0,0.00864)t(A,B,C)=(0.25871,0,-0.00864)\,t when the axes are chosen such that the $k_{x}$ direction is on the $MK$ line and the $k_{y}$ direction on the $M\Gamma$ line. To define the energy dispersion of the Taylor model on the complete hexagonal Brillouin zone, we divide the latter into six equilateral triangles. Each of them has a single $M$ point on its boundary. We define the energy dispersion as a piecewise function on these triangles, where in each triangle the canonical dispersion of the corresponding $M$ point is used. This approach conserves the lattice symmetry and ensures that the dispersion is continuous. It is, however, not differentiable on the $\Gamma K$ lines.
@@ C Low energy effective model
We employ the functional renormalization group in its standard static four-point approximation, neglecting self-energy corrections, frequency dependencies, and all interaction vertices with more than 4 fermionic fields, i.e. two interacting particles [97, 71, 87, 90, 5]. Since self-energy corrections enter only at higher loop order in the renormalization of the scattering vertex, they typically remain subleading and do not affect the qualitative tendencies toward phase transitions considered here [71, 21]. The main methodological improvement compared to previous studies of HOVHS [21, 34, 11, 123, 39] is that within the aforementioned approximations our FRG analysis systematically accounts for all possible ordering tendencies and fluctuations representable as fermionic bilinears with arbitrary momentum transfers $\boldsymbol{q}$ in an unbiased fashion. While secondary momenta $\boldsymbol{k}$ and $\boldsymbol{k}^{\prime}$ are expanded in real-space form factors that are truncated at some distance $\alpha$ (truncated unity FRG formulation), transfer momenta are sampled on equidistant momentum points within the Brillouin zone, where the modern implementation used in the present work [90, 5, 89, 88] allows for high resolutions up to $\approx 25000$ momentum mesh points within the BZ.
@@
Since the FRG is only valid directly at the phase transition—where the divergence is encountered—the mean-field decomposition is performed at the phase boundary where $\Delta\rightarrow 0$. Consequently, we can linearize the self-consistent gap equation, recasting it as an eigenvalue problem for each $\boldsymbol{q}$ directly at the phase transition. The maximum of eigenvalues at each transfer momentum in the particle-hole channel of our model, indicating the leading fluctuation, is displayed on an exemplary momentum mesh in Fig. 2a in the main text. The associated eigenstate of a divergent eigenvalue defines the structure of the leading order parameter.
@@ A Interaction parameter phase diagrams
In general, identifying a suitable set of bare interaction parameters from first principles is challenging. While constrained random phase approximation (cRPA) calculations seem to suggest a universal scaling behavior for long-range interactions in several kagome compounds [22], the present model features several key differences: The kagome bandstructure is constrained to the material’s surface, is fully spin polarized, and exhibits a breathing distortion. Additionally, cRPA calculations are only suitable for predicting the relative strength of different interaction terms and do not allow for a faithful estimation of the absolute interaction scale [36, 92].
To circumvent these problems, we perform FRG interaction parameter scans. We add density-density interactions to the tight-binding Hamiltonian up to fourth nearest neighbor distance:
@@
The predominant part of parameter space of the breathing model consists of the unprecedented $\sqrt{3}\times\sqrt{3}$ density wave order originating from the diffuse nesting mechanism introduced by the present work. In the limit of weak initial interactions, superconductivity and nematic orders emerge as the dominant tendencies. The latter is in agreement with previous theoretical studies of HOVHS scenarios in hexagonal lattice systems [21, 34] and theoretical and experimental studies of Co3Sn2S2 [74]. In experiment, such nematic order may be further stabilized by strain, deviations from the exact VH filling, or as a consequence of weak local correlations at the Co surface—situating the system in the weak coupling region of the phase diagram in which $\boldsymbol{q}=\Gamma$ orders dominate. Upon introducing a spin degree of freedom, the leading instabilities reveals a magnetic ordering tendency rather than charge order; the alternating bond phase then corresponds to the spins rather than charge density. While we comprehensively analyzed the mean field equations and Ginzburg-Landau free energy for the scenario of spinless fermions, we leave a comparably thorough study of magnetic Kekulé order to future works.
@@ B $K$-order stability analysis
To extend the stability analysis of the $K$-ordered phase beyond the interaction parameter regimes, we vary both the kinetic parameters and the chemical potential of the model given in Eq. SM 1 and repeat the FRG analysis. The default parameters, used both in the theory section of Nag et al. [74] and our calculations, are those of the mathematically precise HOVHS.
B.1 Deviations from Van Hove filling
We observe a suppression of the $K$-ordering tendency upon small doping. While the order persists within the regime of intermediate initial interaction strengths for dopings $\Delta$$\mu\leq 0.05t$, it is absent within the interaction parameter range accessible to the FRG calculation for $\Delta$$\mu\geq 0.1t$. A strong susceptibility to changes in ordering tendencies for small deviations from Van Hove filling is a well-known feature of FRG studies. Comparing the suppression of instabilities with finite transfer momentum—induced by detuning the chemical potential—between the diffuse nesting scenario here with standard Van Hove nesting, we find that in the diffuse nesting case a larger chemical potential change is tolerated. The reference model with standard Van Hove nesting is a kagome Hubbard model at the $p$- and $m$-type Van Hove filling, analyzed via FRG calculations for various values of $\Delta\mu$ as performed in Ref. [91]. The strong stability against detuning of the chemical potential, compared to other systems, is explained by its dependence on scattering events within a “diffuse” energy window around the Fermi energy. Since this energy window is extended, a small deviation of chemical potential does not completely remove the overlap.
B.2 Deviations from the higher-order Van Hove scenario
As outlined earlier, a van Hove singularity is classified as higher-order if at least one of the Hessian eigenvalues vanishes. We analyze the effect of small deviations from this scenario by varying the hopping parameters of the model. To analyze the stability of the Kekulé order at the experimentally accessible DFT fit parameters that approximately correspond to Co3Sn2S2 [74] (see Section A), we interpolate between these and the exact HOVHS. This reintroduces a finite quadratic term along the $\Gamma M$ line. At sufficiently flat quadratic dispersion, the $K$-order persists, demonstrating that the flatness-induced high DOS of the VHS close to its higher-order configuration drives the mechanism and an exact vanishing of the Hessian determinant is not required. As the deviation is increased, we observe that the transfer momentum associated with the ordering tendency shifts toward an incommensurate momentum point located along the $KM$ line. We find that this is consistent with a shift in the momentum of the particle-hole susceptibility peak, further substantiating the claim that diffuse nesting is purely of kinetic origin. However, if longer-range interactions that favor the charge imbalance between neighboring up-pointing triangles ($V_{34}$) are strong, the $K$-order can be stabilized at the DFT fit parameters even though the particle-hole susceptibility peak shifts to incommensurate momenta. The stability of the $K$-order towards different hopping terms can be explained by the possibility to construct a universal single-band hexagonal $M$-HOVHS model, and by the diffuse nesting mechanism itself: Since we do not require an exact overlap of Fermi surfaces or of high-DOS points (as in perfect nesting or Van Hove nesting scenarios), the $K$-order is stable to a relatively large uncertainty in the hoppings.
@@ A Overlap of broadened Fermi surfaces
The particle-hole susceptibility peak at transfer-momentum $\boldsymbol{q}=K$ in Fig. 3a (main text) is traced back to the diffuse Fermi surface (FS) nesting inherent to the warped FS emerging in the presence of HOVHSs, where the quartic term points along the $\Gamma M$ direction. The Fermi surface overlaps at $\boldsymbol{q}=K$ and $\boldsymbol{q}=M$, with broadening realized by a finite temperature $T=0.01t$, are shown as heatmaps in Fig. SM 4. The gray regions indicate the two shifted FS. In the $K$-shifted scenario extended circular nesting regions that drive a susceptibility peak appear, while the nesting at $M$-shifting remains poor. This is in contrast to the perfect VHS nesting emerging for ordinary VHS at transfer momenta $M$ in hexagonal lattice systems with only (real) nearest neighbor hybridization.
@@
Figure SM 4: Fermi surface overlaps of breathing and non-breathing kinetic models with nesting-vectors $\boldsymbol{q}=M$ and $\boldsymbol{q}=K$ at HOVHS filling. The broadening factor is T=0.01tT=0.01\,t. Upon $K$-shifting circular regions appear, indicating strong nesting. When shifted by $M$, the nesting is poor.
B Causes of Fermi surface broadening in kinetic and interacting models
It is important to clearly distinguish the FS broadening in the kinetic models, where it is implemented by hand to investigate its effects on susceptibilities, from that in the interacting models of the FRG calculation.
In the kinetic model, the susceptibility in band space (cf. Eq. (4) in the main text) reduces, after evaluation of the Matsubara sum, to the Lindhard formula:
$\chi^{0}_{nm}(\bm{q})=-\frac{1}{\beta}\int_{\mathrm{BZ}}\text{d}\bm{k}\,\frac{f(\epsilon_{n}(\bm{k}+\bm{q}))-f(\epsilon_{m}(\bm{k}))}{\epsilon_{n}(\bm{k}+\bm{q})-\epsilon_{m}(\bm{k})+i\delta^{+}},$ (SM 4)
where $f$ is the Fermi-Dirac distribution and $\epsilon_{n}(\bm{k})$ denotes the dispersion of the $n$-th band. There are multiple ways to broaden the effective energy window of the dominant scattering processes: The first is finite temperature, which modifies the numerator by “smearing out” the Fermi functions. The other is a finite quasiparticle lifetime due to interactions or disorder that introduce an imaginary part to the denominator, modeled with a finite $\delta^{+}$, which results in a Lorentzian broadening. Both variations have the same effect of driving diffuse $K$-nesting by allowing scattering at non-zero energy.
In the FRG calculation for the interacting model, the RG scale $\Lambda$ effectively acts as broadening, playing a similar role as temperature [4, 59, 25]. Even though our approximations for the FRG fix $T=0$, this effectively broadens the Fermi surface and therefore allows for diffuse nesting.
Thermal broadening, which is used throughout the main section, is only one possibility to implement a broadened Fermi surface within the non-interacting models. While therefore a finite temperature is not expected to harm the diffuse nesting mechanism, it is not required, and the mechanism is expected to be stable at zero temperature.